OpenStudy (anonymous):

one side of a square is 3x +2. which expression represents the perimeter

OpenStudy (ddcamp):

A square has 4 sides of equal length, right? So, the perimeter would be 4×(The length of one side) = 4×(3x+2) = 4(3x) + 4(2) = 12x + 8